OpenStudy (anonymous):

Use implicit differentiation to find dy/dx cosxy + x^7 = y^7

OpenStudy (ash2326):

We have \[\cos xy+x^7=y^7\] we know the differentiation of the following \[\frac{d}{dx} cos x=-sin x\] and \[\frac{d}{dx} x^n=nx^{n-1}\] Let's use these here \[\frac{d}{dx}(\cos xy +x^7)=\frac{d}{dx} y^7\] we get \[-\sin xy \times \frac{d}{dx}(xy)+7x^6=7y^6 \frac{dy}{dx}\] \[-\sin xy (y+x\frac{dy}{dx})+7x^6=7y^6 \frac{dy}{dx}\] we get \[-y \sin xy-x \sin xy \frac{dy}{dx}+7x^6=7y^6 \frac{dy}{dx}\] or \[-y \sin xy- 7x^6=7y^6 \frac{dy}{dx}+ \sin xy \frac{dy}{dx}\] Or we get \[\frac{dy}{dx}=\frac{-y \sin xy- 7x^6}{\sin xy +7y^6}\]